OpenStudy (anonymous):

Let \(X, Y, Z\) be metric spaces. \(f: X \rightarrow Y\) is called uniformly continuous if \(\forall \epsilon > 0 \quad \exists \delta > 0 \quad \forall x, y \in X : d_{x}(x,y) \leq \delta \Rightarrow d_{Y}(f(x),f(y))\leq \epsilon \) Show: b) \(f(x):=x^{2}\) is continuous, but not uniformly continuous on \(\mathbb{R}\)

OpenStudy (anonymous):

you can proove the not uniformly continuous part by finding counterexample. Try to take: \[\epsilon=1\] and work out the rest from definition.

OpenStudy (anonymous):

myko thank you very much, but i need complete solution because about 40 min ago i should it deliver to my mentor... i mean i dont have time try to understand it and solve....

OpenStudy (anonymous):

any help would be apreciated

OpenStudy (anonymous):

then just look maybe here: http://answers.yahoo.com/question/index?qid=20101103061142AAP8pTE

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

there it says\[[1,\infty]\] but we have in question about \[\mathbb{R}\] is it much difference?

OpenStudy (anonymous):

If it isn't uniformly continuous on a subset of \(\mathbb R\) then it certainly isn't uniformly continuous on \(\mathbb R\).

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Take \[ x_n = n + \frac 1 n\\ y_n= n - \frac 1 n\\ x_n - y_n= \frac 2n\\ x_n^2 - y_n^2= 4\\ \lim_{n\to \infty} x_n -y_n=0 \\ \text { but } \lim_{n\to \infty} x_n^2 -y_n^2=4 \] So f(x) is not uniformly continuous.